# PrivyWave: Privacy-Aware Wireless Sensing of Heart Beat

### v1

@@ 6.3. Wearability and Form Factor
Our current prototype requires connection to an external power source due to the pneumatic actuators’ power demands, limiting mobility to stationary monitoring scenarios. Future work should focus on miniaturization and power optimization to enable battery-powered operation. Beyond power considerations, seamless integration into everyday clothing would significantly improve usability. Recent advances in smart textiles provide promising directions (Honnet et al., 2025): actuators and circuits could be embedded within fabric layers near the chest area, or shirt buttons could be redesigned as miniature pneumatic actuators. Such integration would eliminate the need for dedicated wearable devices while maintaining obfuscation effectiveness without compromising comfort or aesthetics.
@@ 6.4. Other Vital Signs
While this work focused on heartbeat protection, wireless sensing can detect other involuntary physiological signals with privacy implications. Breathing generates larger movements (millimeters vs. sub-millimeter) at lower frequencies (0.1-0.5 Hz), making it more detectable but potentially easier to obfuscate with larger-displacement actuators. Beyond cardiorespiratory signals, wireless sensors can detect other health related informatoin such as tremors (Lo et al., 2011)(Parkinson’s disease, 4-12 Hz). This will revealing private health conditions that could lead to discrimination. Future work could extend obfuscation techniques for breathing and other health related information leakage.
@@ 7. Conclusion
We presented PrivyWave, a key-based physical obfuscation system for selective privacy protection in wireless heartbeat sensing. By generating controlled decoy heartbeat signals at cryptographically-determined frequencies, our system enables authorized sensors to recover accurate measurements while unauthorized sensors cannot distinguish true signals from decoys. Our evaluation across mmWave radar and acoustic sensing demonstrates effective protection (average unauthorized error: 21.3-42.0 BPM) while maintaining high authorized accuracy (5.8-9.7 BPM). The authorized accuracy is comparable to typical wireless sensing heart rate measurement systems (ranging from 3-15 BPM) (Qian et al., 2018; Alizadeh et al., 2019), meaning PrivyWave does not hamper the utility of wireless sensing. The system operates across multiple sensing modalities without per-modality customization and provides formal security guarantees through cryptographic key-based decoding. This work establishes physical-layer obfuscation as a viable approach for balancing privacy and utility in pervasive health monitoring, opening new directions for privacy-preserving sensing systems.

### v2

@@ 6.3. Wearability and Form Factor
Our current prototype requires connection to an external power source due to the pneumatic actuators’ power demands, limiting mobility to stationary monitoring scenarios. Future work should focus on miniaturization and power optimization to enable battery-powered operation. Beyond power considerations, seamless integration into everyday clothing would significantly improve usability. Recent advances in smart textiles provide promising directions (honnet2025fibercircuits): actuators and circuits could be embedded within fabric layers near the chest area, or shirt buttons could be redesigned as miniature pneumatic actuators. Such integration would eliminate the need for dedicated wearable devices while maintaining obfuscation effectiveness without compromising comfort or aesthetics.
@@ 6.4. Other Vital Signs
While this work focused on heartbeat protection, wireless sensing can detect other involuntary physiological signals with privacy implications. Breathing generates larger movements (millimeters vs. sub-millimeter) at lower frequencies (0.1-0.5 Hz), making it more detectable but potentially easier to obfuscate with larger-displacement actuators. Beyond cardiorespiratory signals, wireless sensors can detect other health related informatoin such as tremors (lo2011wireless)(Parkinson’s disease, 4-12 Hz). This will revealing private health conditions that could lead to discrimination. Future work could extend obfuscation techniques for breathing and other health related information leakage.
@@ 7. Conclusion
We presented PrivyWave, a key-based physical obfuscation system for selective privacy protection in wireless heartbeat sensing. By generating controlled decoy heartbeat signals at cryptographically-determined frequencies, our system enables authorized sensors to recover accurate measurements while unauthorized sensors cannot distinguish true signals from decoys. Our evaluation across mmWave radar and acoustic sensing demonstrates effective protection (average unauthorized error: 21.3-42.0 BPM) while maintaining high authorized accuracy (5.8-9.7 BPM). The authorized accuracy is comparable to typical wireless sensing heart rate measurement systems (ranging from 3-15 BPM) (qian2018acousticcardiogram; alizadeh2019remote), meaning PrivyWave does not hamper the utility of wireless sensing. The system operates across multiple sensing modalities without per-modality customization and provides formal security guarantees through cryptographic key-based decoding. This work establishes physical-layer obfuscation as a viable approach for balancing privacy and utility in pervasive health monitoring, opening new directions for privacy-preserving sensing systems.